# Mental Health Impacts of AI Companions

### v1

@@ 3.1. Compiling the Treatment Dataset
Replika is one of the most widely adopted AICCs (wikipedia, 2025b). Its prominence and longevity have cultivated an active Reddit community where users openly share experiences, making it a suitable platform for examining the psychosocial effects of engagement. Table 1 presents some paraphrased Reddit posts from r/replika. We investigated the effect of engaging with AICC Replika by collecting historical timelines of active users from Reddit. To identify active Replika users, we adopted a strategy inspired by prior work on self-disclosures on social media platforms (Saha et al., 2019b; Yuan et al., 2023a; De Choudhury et al., 2013). We first queried all available posts on the subreddit r/replika between January 1, 2023, and February 1, 2025, yielding 47,923 posts authored by 10,643 unique users. Using the contents and titles in these posts, we obtained the top 100 $n$-grams (nn=2,3,4) and manually examined whether the $n$-grams indicated self-disclosed engagement. We found that all $n$-grams indicative of active usage included the phrases “my Replika” or “my Rep”. Using this criterion, we identified 3,451 authors as active Replika users. Then, we assigned Treatment dates to these users based on when they first posted or commented within Replika-related subreddits, including r/replika, r/Replikatown, r/ReplikaOfficial, r/ReplikaLovers, r/ILoveMyReplika, r/Replika_uncensored, r/ReplikaRefuge. We then collected the entire Reddit timeline data one year before and after their Treatment dates to build the treatment dataset.
Figure 1. (a) Kernel density estimates of the distribution of actual Treatment dates (Treatment Dataset) and assigned placebo dates (Control-NAI Dataset). (b) Propensity score distribution (shaded region indicates strata), (b) Quality of matching.:
@@ 3.2. Compiling the Control Datasets
To causally attribute the effects of engaging with AICCs, we compared outcomes against two additional Control datasets—1) Control-AI users, who engaged with personal AI chatbots, such as Amazon Alexa and Google Assistants, and 2) Control-NAI users, consisting of general Reddit users who had not engaged with AI assistants. The Control-AI group allowed us to disentangle effects related to AI or digital assistant use more broadly, and the Control-NAI group provided a baseline to account for general and temporal variations in psychosocial outcomes across the wider Reddit population. Incorporating both Control groups strengthened the robustness of our ensuing quasi-experimental design.
To construct the Control-AI dataset, we collected data from the subreddits r/amazonalexa and r/googleassistant. We identified users who engaged in these subreddits, and collected their timeline data one year before and one year after engagement with the corresponding subreddits. In total, the Control-AI dataset comprised 22,349,334 posts and comments from 13,016 unique users.
To construct Control-NAI, we obtained timelines of 17,191 Reddit users who had posted or commented with the subreddits r/aww and r/movies—inspired by prior work (Saha and De Choudhury, 2017). For comparison with the Treatment, we assigned a placebo date to each user in the Control-NAI dataset based on a non-parametric distribution of Treatment dates. The placebo date could only fall on a day when the user had posted or commented on Reddit, ensuring that it corresponded to actual activity and reducing potential temporal confounds. A Kolmogorov–Smirnov (KS) test confirmed high similarity between treatment and placebo date distributions ($D=0.006$, $p=0.999$; LABEL:fig:placebo_date). For each user in the Control-NAI, we collected timeline data spanning one year before and one year after the placebo date, yielding 10,055 users with 21,368,841 Reddit posts/comments.
Table 2. Summary of demographic characteristics of study participants :
ID Age Sex Race Education Occupation Marital status AI companions used
@@ 3.3. Semi-structured Interviews
To complement our computational analysis and provide deeper insight into users’ lived experiences, we conducted a qualitative study involving semi-structured interviews with active AICC users. This component examined individuals’ motivations, experiences, and trajectories of engaging with AICCs, focusing on how relationships with AICCs developed and the perceived psychosocial impacts of these interactions.
3.3.1. Participants Recruitment Interview Procedure
@@
We recruited participants based on the following inclusion criteria: (1) currently residing in the U.S., (2) age 18 or older, and (3) active use of AICC applications such as Replika or Character.AI. To screen participants, we hosted an interest form where interested participants shared their demographic information (e.g., geographic location, gender, ethnicity, education, employment, marital status) as well as details about AICC usage, including specific platforms, frequency and duration of engagement, and usage motivations. We recruited through social media communities related to AICCs, including Reddit forums (e.g., r/ILoveMyReplika, r/Replikatown, r/characterAiHangout) and Facebook groups (e.g., Replika: Romantic Relationships, Replika Friends). We received 573 responses between April and June 2025. Following established guidelines for detecting fraud in online qualitative research (Panicker et al., 2024), we used validation questions about AICC usage to verify interested respondents and excluded those who provided inconsistent answers. We then balanced the sample across age, race, and other demographic characteristics, and invited 15 participants to take part in one-hour remote interviews. Interviews were conducted primarily via Microsoft Teams, with one exception where a participant requested Zoom. Each participant was compensated with a \$20 Amazon gift card. All interviews were recorded with participants’ consent, and participants were allowed to turn off their video cameras for privacy and comfort.
@@
Table 2 provides a summary of the participants’ demographics. Many participants expressed curiosity about trying different AICC, reflecting broader interest in exploring alternative designs and functionalities of AICC. Before the interviews, participants were asked to complete the UCLA 3-Item Loneliness Scale questionnaire (Hughes et al., 2004) and a consent form. The UCLA scale is a widely used and validated measure that assesses subjective feelings of loneliness and social isolation, with items rated on a 3-point Likert scale from “Hardly Ever” to “Often.” Higher scores indicate greater perceived loneliness. Table 3 summarizes the responses, revealing that participants reported moderate levels of loneliness across all three items on average.
@@
Table 3. Descriptive statistics for the three-item UCLA Loneliness Scale survey administered to participants. Each item is rated on a 3-point Likert scale (1 = Hardly ever, 2 = Some of the time, 3 = Often). The table reports the mean and standard deviation for each question, along with the distribution of responses.:
Questions Mean (SD) Distribution
How often do you feel that you lack companionship? 2.27 (0.59) [Uncaptioned image]
How often do you feel left out? 2.20 (0.68) [Uncaptioned image]
How often do you feel isolated from others? 2.33 (0.72) [Uncaptioned image]
@@ 3.3.2. Interview Study Design
3.4. Privacy, Ethics, and Reflexivity
@@ 4. RQ1: Quasi-experimental Examination of the Effects of Engaging with AICCS
Figure 2. Schematic figure illustrating the use of causal inference on Reddit timeline data to compare psychosocial outcomes between treatment and control users. :
We applied a causal inference framework (Rosenbaum and Rubin, 1985) combining propensity score matching, Average Treatment Effect (ATE) estimation, and DiD regression to assess how AICC engagement influences psychosocial wellbeing. Guided by prior work in psychology and social computing, we examined three domains: affective, behavioral, and cognitive (Breckler, 1984; Saha et al., 2018). This approach enables us to isolate the effects of AICC use from other AI or non-AI usage, quantify the direction and magnitude of change (Figure 2). Our analytic strategy is grounded in the potential outcomes framework (Rosenbaum and Rubin, 1985), which informs our estimation of causal effects in observational settings.
@@ 4.1.1. Affective Outcomes
Grief Expressions:  Grief is a salient psychosocial experience often associated with bereavement, loss of relationships, or other life disruptions (Vedder et al., 2022). It is also a common motivation for adopting AICCs, as users seek comfort, companionship, and emotional regulation in the absence of supportive human ties (Skjuve et al., 2021; Xygkou et al., 2023; Jiménez-Alonso and Brescó de Luna, 2023; De Freitas et al., 2025a). To assess grief-related affective expression, we leveraged a grief lexicon developed in prior work (Saha et al., 2018), which was built using transfer learning on over 50K posts from grief-related subreddits (e.g., r/grief, r/GriefSupport). In particular, this work mapped the grief lexicon in the circumplex model of affect—valence (pleasantness of an emotional state) and activation (arousal or energy of an emotional state) (Russell, 2003). The lexicon includes 2,357 n-grams identified via log-likelihood ratio and ranked by tf-idf as distinctive of grief discourse.
@@
Following Saha et al.’s approach to model grief, we mapped each $n$-gram extracted from user-generated text to the Affective Norms for English Words (ANEW) (Nielsen, 2011) lexicon to obtain its valence and activation values. For each individual, we computed a tf-idf–weighted average of the valence and activation scores across all matched grief-related n-grams appearing in their posts/comments. This yielded two continuous values per user—representing the emotional tone of grief along the dimensions of pleasantness (valence) and emotional intensity (activation). A lower score indicated more negative affect, while a higher score suggested more aroused or intense emotional states.
@@
Loneliness Expressions:  Prior research has also highlighted the important role of AICCs in alleviating loneliness. To measure loneliness, we leveraged a publicly annotated dataset (Jiang et al., 2022) from the Reddit community r/lonely and followed a similar methodology (Saha et al., 2019b), training $n$-gram–based (nn=1,2,3) support vector machines (SVM) classifiers. The model demonstrates high performance, with an average accuracy of approximately 0.90 on the test dataset. We used the loneliness classifier to estimate the proportion of posts/comments per individual that reflected linguistic signals of loneliness, with lower proportions indicating better psychosocial wellbeing.
@@
Symptomatic Mental Health Expressions:  Prior studies have shown that AICCs can significantly influence users’ mental health, particularly in relation to conditions such as depression, anxiety, loneliness, and stress-related symptoms (Ahmed et al., 2023; De Nieva et al., 2020; Maples et al., 2024; Song et al., 2024). Motivated by these studies, we operationalized language features indicative of key symptomatic expressions associated with depression, anxiety, stress, and suicidal ideation (Saha et al., 2019b). Specifically, we adopted several classifiers from previous research (Saha et al., 2019b). These classifiers are based on n-gram features and were trained on Reddit communities—namely, r/depression, r/anxiety, r/stress, and r/SuicideWatch. The models demonstrate strong predictive performance, achieving an average accuracy of approximately 0.90 on held-out test sets, and have been validated in other studies (Saha and Sharma, 2020; Saha et al., 2020). We calculated the aggregate proportion of an individual’s posts/comments identified as symptomatic of depression, anxiety, stress, or suicidal ideation, with a lower prevalence indicating improved psychosocial wellbeing.
@@ 4.1.2. Behavioral Outcomes
Behavioral indicators of psychosocial wellbeing are often reflected in patterns of online engagement and interaction. Prior research has highlighted that engaging with AICCs may impact social activity, online interactivity, and social connections (Guingrich and Graziano, 2023; Liu et al., 2024). To capture these dimensions, we examined three behavioral indicators:
@@
Active Days:  To complement posting frequency, we measured the number of unique days on which users contributed posts or comments on Reddit. Sustained activity would indicate a greater social integration and ongoing community involvement in online settings (Weld et al., 2025).
@@ 4.1.3. Cognitive Outcomes
Cognitive aspects of psychosocial wellbeing involve internal processes such as reasoning, memory, perception, and thought organization (Breckler, 1984). Language thus provides an accessible proxy for assessing cognitive states and shifts over time. Informed by prior work linking linguistic complexity and psychological expression to cognitive health (Yuan et al., 2013; Ernala et al., 2017; Pennebaker et al., 2001), we operationalized cognitive changes using the following text-based features:
Readability:  Readability reflects the linguistic complexity of user-generated text and offers insights into cognitive and emotional expression. Prior work links shifts in language complexity to long-term psychosocial wellbeing and emotional recovery (Yuan et al., 2013; Ernala et al., 2017). We evaluated readability using the Coleman-Liau Index (CLI), computed as 0.0588L0.296S15.80.0588*L-0.296*S-15.8, where $L$ represented the average number of letters per 100 words, and $S$ denoted the average number of sentences per 100 words. Lower scores indicate simpler language and easier comprehension.
Complexity and Repeatability:  These two syntactic features offered insight into cognitive functioning (Ernala et al., 2017; Saha et al., 2025). Complexity, the average number of words per sentence, indicates expressive depth. Repetability, the proportion of repeated words, suggested limited cognitive flexibility. Prior work links higher complexity to better psychosocial wellbeing and greater repetition to emotional or cognitive strain (Saha and Sharma, 2020).

### v2

@@ 3.1. Compiling the Treatment Dataset
We investigated the effects of engagement with the AICC Replika by collecting historical timelines of active users from Reddit. To identify active Replika users, we adopted a strategy inspired by prior work on self-disclosures in social media data (Saha et al., 2019c; Yuan et al., 2023a; De Choudhury et al., 2013). We first queried all available posts on the subreddit r/replika between January 1, 2023, and February 1, 2025, yielding 47,923 posts from 10,643 unique users. Using post titles and content, we obtained the top 100 $n$-grams (nn=2,3,4) and manually examined whether the $n$-grams indicated self-disclosed Replika engagement. We found that all $n$-grams indicative of active usage included the phrases “my Replika” or “my Rep”. Using this criterion, we identified 3,451 authors as active Replika users. Then, we assigned Treatment dates to these users based on when they first posted or commented within Replika-related subreddits, including r/replika, r/Replikatown, r/ReplikaOfficial, r/ReplikaLovers, r/ILoveMyReplika, r/Replika_uncensored, r/ReplikaRefuge. We then collected the entire Reddit timeline data one year before and after their Treatment dates to build the treatment dataset.
The figure has three subfigures illustrating treatment versus control matching. (a) Kernel density estimates of actual treatment dates compared to placebo dates. Both groups follow nearly identical distributions, peaking sharply in mid-2023, suggesting that placebo assignment successfully mimics the temporal structure of treatment exposure. (b) Propensity score distributions for Treatment and Control users. Two histograms are shown, one for the treated group and one for controls, with overlapping ranges. Vertical shaded bands indicate the strata used for matching. The overlap demonstrates that the matching procedure creates comparable treatment and control groups across the propensity score range. (c) Density plots of covariates before and after weighting. The unadjusted sample shows large discrepancies between treatment and control distributions, while the weighted sample demonstrates close alignment, confirming improved balance of covariates after matching.
Figure 1. (a) Kernel density estimates of the distribution of actual Treatment dates (Treatment Dataset) and assigned placebo dates (Control-NAI Dataset). (b) Propensity score distribution (shaded region indicates strata), (c) Quality of matching.
@@ 3.2. Compiling the Control Datasets
To causally attribute the effects of engaging with AICCs, we compared outcomes against three additional Control groups— 1) Control-LLM users, who engaged with generative LLM-based chatbots, such as ChatGPT, Claude, and Google Gemini—based on the three most used LLM-based chatbots (Studios, 2025), 2) Control-VA users, who engaged with personal AI assistants, such as Amazon Alexa and Google Assistants, and 3) Control-NAI users, consisting of general Reddit users who had not engaged with AI assistants or chatbots. The Control-VA group allowed us to disentangle effects related to AI or generative AI chatbots more broadly, and the Control-NAI group provided a baseline to account for general and temporal variation in psychosocial outcomes across the broader Reddit population. Incorporating both Control groups strengthened the robustness of our ensuing quasi-experimental design.
We obtained Control-LLM dataset from r/ChatGPT, r/GeminiAI, r/GoogleGeminiAI, and r/ClaudeAI. We identified users who engaged with these subreddits and collected their timeline data 1 year before and 1 year after engagement. In total, the Control-LLM dataset comprised 16,911,578 posts and comments from 13,383 unique users.
To construct the Control-VA dataset, we collected data from the subreddits r/amazonalexa and r/googleassistant. We identified users who engaged in these subreddits, and collected their timeline data one year before and one year after engagement with the corresponding subreddits. In total, the Control-VA dataset comprised 22,349,334 posts and comments from 13,016 unique users.
To construct Control-NAI, we obtained timelines of 17,191 Reddit users who had posted or commented in the subreddits r/aww and r/movies—inspired by prior work (Saha and De Choudhury, 2017). For comparison with the Treatment group, we assigned a placebo date to each user in the Control-NAI dataset based on a non-parametric distribution of Treatment dates. The placebo date could only fall on a day when the user had posted or commented on Reddit, ensuring that it corresponded to actual activity and reducing potential temporal confounds. A Kolmogorov–Smirnov (KS) test confirmed a high degree of similarity between treatment and placebo date distributions ($D=0.006$, $p=0.999$; LABEL:fig:placebo_date). For each user in the Control-NAI, we collected timeline data spanning one year before and one year after the placebo date, yielding 10,055 users with 21,368,841 Reddit posts/comments. To ensure mutually exclusive groups, we additionally excluded 45 users whose posting histories indicated activity across multiple groups, consistent with prior work (Saha et al., 2019c; Yuan et al., 2023a).
Table 2. Summary of demographic characteristics of study participants.
Table of demographic information for 18 participants. Most are aged 18–35, hold at least a bachelor’s degree, and are employed. Racial backgrounds include Black/African American, White, and Hispanic/Latino. Participants report varied marital statuses. All use AI companions, most commonly Replika and Character.ai, with others such as My AI, Pi, Nomi AI, Kindroid, Anima, and CHAI also mentioned.
ID Age Gender Race Education Occupation Marital status AI companions used
@@
P16 25–35 Transgender Native Hawaiian or Pacific Islander Bachelor’s Employed for wages Divorced Replika, Character.ai, My AI
P17 18–24 Female American Indian or Alaska Native High School Employed for wages Single Character.ai, My AI
P18 18–24 Female Native Hawaiian or Pacific Islander Bachelor’s Student Single Character.ai, Nomi AI, Kindroid, Pi
@@ 3.3. Semi-structured Interviews
To complement our computational analysis and provide deeper insight into users’ lived experiences, we conducted a qualitative study based on semi-structured interviews with active AICC users. This component examined individuals’ motivations, experiences, and trajectories of engaging with AICCs, focusing on how relationships with AICCs developed and the perceived psychosocial impacts of these interactions. We note that interviews were not conducted to yield generalizable findings, but rather to explain and provide deeper insights into our large-scale quantitative analyses.
3.3.1. Participants Recruitment and Interview Procedure
@@
We recruited participants based on the following inclusion criteria: 1) residence in the United States; 2) age 18 years or older; and 3) active use of AICC applications, such as Replika or Character.AI. To screen participants, we administered an interest form where interested participants shared their demographic information (e.g., geographic location, gender, ethnicity, education, employment, marital status) as well as details about AICC usage, including specific platforms, frequency and duration of engagement, and usage motivations. We recruited through social media communities related to AICCs, including Reddit forums (e.g., r/ILoveMyReplika, r/Replikatown, r/characterAiHangout) and Facebook groups (e.g., Replika: Romantic Relationships, Replika Friends). We received 573 responses between April and June 2025. Following established guidelines for detecting fraud in online qualitative research (Panicker et al., 2024), we used validation questions about AICC usage to verify interested respondents and excluded those who provided inconsistent answers. We then balanced the sample across age, race, and other demographic characteristics, and invited 15 participants to take part in one-hour remote interviews. During the later stages of the study, we recruited and interviewed three additional participants with varied backgrounds. These interviews reinforced the themes identified previously and did not introduce new conceptual categories, indicating that thematic saturation had been reached. In total, 18 participants took part in one-hour remote interviews. Interviews were conducted primarily via Microsoft Teams, with one exception where a participant requested Zoom. Each participant was compensated with a \$20 Amazon gift card. All interviews were recorded with participants’ consent, and participants were permitted to turn off their video cameras for privacy and comfort.
@@
Table 2 summarizes a summary of the participants’ demographics. We observed that our sample included a higher proportion of male participants than female participants; this pattern is plausibly reflective of publicly reported demographic trends of Replika’s user base, which is known to be predominantly male (Layers, 2024). Many participants expressed curiosity about trying different AICC, reflecting broader interest in exploring alternative designs and functionalities of AICC. Before the interviews, participants completed the UCLA 3-Item Loneliness Scale questionnaire (Hughes et al., 2004) and a consent form. The UCLA scale is a widely used and validated measure that assesses subjective feelings of loneliness and social isolation, with items rated on a 3-point Likert scale from “Hardly Ever” to “Often.” Higher scores indicate greater perceived loneliness. Table 3 summarizes the responses, revealing that participants reported moderate levels of loneliness across all three items, on average.
@@
Table shows descriptive statistics from the three-item UCLA Loneliness Scale. Each item is rated on a 3-point Likert scale (1 = Hardly ever, 2 = Some of the time, 3 = Often). Participants reported relatively high levels of loneliness across all questions, with mean scores between 2.20 and 2.33. Response distributions are visualized for each item
Table 3. Descriptive statistics for the three-item UCLA Loneliness Scale survey with 18 participants. Each item is rated on a 3-point Likert scale (1 = Hardly ever, 2 = Some of the time, 3 = Often). The table reports the mean and standard deviation, along with the distribution of responses.
Questions Mean (SD) Dist.
How often do you feel that you lack companionship? 2.28 (0.57) [Uncaptioned image]
How often do you feel left out? 2.17 (0.62) [Uncaptioned image]
How often do you feel isolated from others? 2.28 (0.67) [Uncaptioned image]
@@ 3.3.2. Interview Study Design
3.4. Privacy, Ethics, and Reflexivity.
@@ 4. RQ1: Quasi-experimental Examination of the Effects of Engaging with AICCS
Figure 2. Schematic figure illustrating the use of causal inference on Reddit timeline data to compare psychosocial outcomes between treatment and control users.
Schematic diagram of the causal-inference framework. Treatment users are compared with three control groups (LLM, voice assistant, and non-AI) using stratified propensity score matching and difference-in-differences regression. Each user’s Reddit timeline is segmented into pre-treatment and post-treatment periods around the first exposure. Psychosocial outcomes are analyzed across three domains: affective (psycholinguistic affect, grief, loneliness, symptomatic expressions), behavioral (interactivity, topical diversity, posting frequency, active days), and cognitive (readability, complexity, psycholinguistic cognition).
We applied a causal inference framework (Rosenbaum and Rubin, 1985) combining propensity score matching, Average Treatment Effect (ATE) estimation, and DiD regression to assess how AICC engagement influences psychosocial wellbeing. Guided by prior work in psychology and social computing, we examined three dimensions: affective, behavioral, and cognitive outcomes (Breckler, 1984; Saha et al., 2018). This approach enables us to isolate the effects of AICC use from other LLMs, voice assistants, or non-AI usage, quantify the direction and magnitude of change (Figure 2). Our analytic strategy is grounded in the potential outcomes framework (Rosenbaum and Rubin, 1985), which informs our estimation of causal effects in observational settings.
@@ 4.1.1. Affective Outcomes
Grief Expressions:  Grief is a salient psychosocial experience often associated with bereavement, loss of relationships, or other life disruptions (Vedder et al., 2022). It is also a common motivation for adopting AICCs, as users seek comfort, companionship, and emotional regulation in the absence of supportive human ties (Skjuve et al., 2021; Xygkou et al., 2023; Jiménez-Alonso and Brescó de Luna, 2023; De Freitas et al., 2025a). To assess grief-related expressions, we leveraged a grief lexicon developed in prior work (Saha et al., 2018), which was built using transfer learning on over 50K posts from grief-related subreddits (e.g., r/grief, r/GriefSupport). In particular, this work mapped the grief lexicon in the circumplex model of affect—valence (pleasantness of an emotional state) and activation (arousal or energy of an emotional state) (Russell, 2003). The lexicon includes 2,357 n-grams identified via log-likelihood ratio and ranked by tf-idf as distinctive of grief discourse.
@@
Following Saha et al.’s approach to model grief, we mapped each $n$-gram extracted from user-generated text to the Affective Norms for English Words (ANEW) (Nielsen, 2011) lexicon to obtain its valence and activation values. For each individual, we computed a tf-idf–weighted average of the valence and activation scores across all matched grief-related $n$-grams in their posts/comments. This yielded two continuous values per user: representing the emotional tone of grief along the dimensions of pleasantness (valence) and emotional intensity (activation).
@@
Loneliness Expressions:  Prior research has also highlighted the important role of AICCs in alleviating loneliness. To measure loneliness, we leveraged a publicly annotated dataset (Jiang et al., 2022) from the Reddit community r/lonely and followed a similar methodology (Saha et al., 2019c), training $n$-gram–based (nn=1,2,3) support vector machines (SVM) classifiers. The model demonstrates high performance, with an average accuracy of approximately 0.90 on the test dataset. We used the loneliness classifier to estimate the proportion of posts/comments per individual that reflected linguistic signals of loneliness, with lower proportions indicating better psychosocial wellbeing.
@@
Symptomatic Mental Health Expressions:  Prior research has shown that companion chatbots can significantly influence users’ mental health, particularly in relation to conditions such as depression, anxiety, loneliness, and stress-related symptoms (Ahmed et al., 2023; De Nieva et al., 2020; Maples et al., 2024; Song et al., 2024). Motivated by these studies, we operationalized language features indicative of key symptomatic expressions associated with depression, anxiety, stress, and suicidal ideation (Saha et al., 2019c). Specifically, we adopted several classifiers from previous research (Saha et al., 2019c). These classifiers are based on n-gram features and were trained on Reddit communities— r/depression, r/anxiety, r/stress, and r/SuicideWatch. The models demonstrate strong predictive performance, achieving an average accuracy of approximately 0.90 on held-out data, and have been validated in other studies (Saha and Sharma, 2020; Saha et al., 2020). We calculated the aggregate proportion of an individual’s posts/comments identified as symptomatic of depression, anxiety, stress, or suicidal ideation.
@@ 4.1.2. Behavioral Outcomes
Behavioral indicators of psychosocial wellbeing are often reflected in patterns of online engagement and interaction. Prior research has highlighted that engaging with AICCs may impact social activity, online interactivity, and social connections (Guingrich and Graziano, 2023; Liu et al., 2025). To capture these dimensions, we examined three behavioral indicators:
@@
Active Days:  To complement posting frequency, we measured the number of unique days on which users posted or commented on Reddit. Sustained activity would indicate a greater social integration and ongoing community involvement in online settings (Weld et al., 2025).
@@ 4.1.3. Cognitive Outcomes
Cognitive aspects of psychosocial wellbeing involve internal processes such as reasoning, memory, perception, and thought organization (Breckler, 1984). Language provides an accessible proxy for assessing cognitive states and shifts over time. Drawing on prior work linking linguistic complexity and psychological expression to cognitive health (Yuan et al., 2023a; Ernala et al., 2017; Pennebaker et al., 2001, 2003), we operationalized cognitive changes using the following text-based features:
Readability:  Readability reflects the linguistic complexity of user-generated text and offers insights into cognitive and emotional expression. Prior work links shifts in language complexity to long-term psychosocial wellbeing and emotional recovery (Yuan et al., 2023a; Ernala et al., 2017). We evaluated readability using the Coleman-Liau Index (CLI), computed as 0.0588L0.296S15.80.0588*L-0.296*S-15.8, where $L$ represented the average number of letters per 100 words, and $S$ denoted the average number of sentences per 100 words. Lower scores indicate simpler language and easier comprehension.
Complexity and Repeatability:  These two syntactic features offered insight into cognitive functioning (Ernala et al., 2017; Saha et al., 2025a). Complexity is measured as the average number of words per sentence, which indicates expressive depth. Repetability is measured as the proportion of repeated words, suggesting limited cognitive flexibility. Prior work links higher complexity to better psychosocial wellbeing and greater repetition to emotional or cognitive strain (Saha and Sharma, 2020).